# Heisenberg spin chain models for realising quantum battery with the aid of Dzyaloshinskii–Moriya interaction

### v1

@@ III.1 Spin chain configurations
III.1.1 Open boundary spin chain configuration
In the open boundary configuration, the spins are arranged linearly under open boundary conditions, such that each spin interacts only with its nearest neighbour [22]. The interaction Hamiltonian is given by
$H_{int}$ $=J\hbar\sum_{i=1}^{n-1}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big{]}$ (11)
@@ III.1.1 Open boundary spin chain configuration
III.1.2 Closed boundary spin chain configuration
In the closed boundary configuration, the last spin interacts with the first spin, forming a closed loop. This implements periodic boundary conditions, ensuring that each spin has two neighbours, even at the boundaries [22]. The interaction Hamiltonian is therefore similar to the open boundary case but extended to include the periodic connection between the first and last spins such that
$H_{int}$ $=J\hbar\sum_{i=1}^{n}\Big{[}(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big{]}$ (12)
@@ III.1.2 Closed boundary spin chain configuration
III.1.3 Supercube spin chain configuration
The supercube configuration is inspired by the connectivity structure of quantum neural networks [50]. In this model, the system qubits are arranged in a one-dimensional array, but their interactions are mapped onto the structure of a cube, where the couplings between qubits are represented as edges. To further enrich the connectivity, additional couplings are introduced along the face diagonals and body diagonals. This extended interaction network allows us to study the effect of higher connectivity on the charging performance of the quantum battery. Consequently, the interaction Hamiltonian becomes more complex, incorporating nearest-neighbour, next-nearest-neighbour, and longer-range couplings, each accompanied by corresponding Dzyaloshinskii–Moriya (DM) interaction terms. The corresponding interaction Hamiltonian is
$H_{int}$ $=\Big{\{}J\hbar\sum_{i=1,i\in Z_{odd}}^{8}[(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}$ (13)
@@ III.1.3 Supercube spin chain configuration
$+\Big{\{}J\hbar\sum_{i=1,2,5,6}[(1+\delta)\sigma_{i}^{x}\sigma_{i+2}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+2}^{y}$
@@
$+\Big{\{}J\hbar\sum_{i=1}^{4}[(1+\delta)\sigma_{i}^{x}\sigma_{i+4}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+4}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+4}^{z}]$
@@
III.2 QB Performance Metrics
@@
III.2.1 Ergotropy
Ergotropy is defined as the maximum amount of work that can be extracted from a quantum battery [51, 52, 53, 54]. The concept relies on the notion of a passive state, from which no work can be extracted under unitary operations [55, 56, 57, 58]. For pure states, the ground state of the system is passive [59]. Accordingly, the ergotropy is given by the difference in energy between the evolved state $|\psi(t)\rangle=U|\psi(0)\rangle$ and the ground state $|\psi(0)\rangle$, i.e.
$\zeta=\text{Tr}(\rho H_{B})-\max{\text{Tr}(U\rho U^{\dagger}H_{B})},$ (14)
where, $\rho$ is the density matrix of the system and $H_{B}$ denotes the internal Hamiltonian of the battery as defined in Eq. (5).
III.2.2 Charging Power
@@
IV Results
In this section we present the results for the eight-qubit quantum battery under different spin-chain configurations and Hamiltonian models. Our analysis focuses on the Ising and XXZ models, and compares their performance in terms of ergotropy and charging power. We also examine the influence of the battery Hamiltonian $H_{B}$ on the charging process by varying the parameter $\lambda\in[0,1]$.
@@ IV Results
Unless stated otherwise, the values of the parameters used in this study are as follows: number of qubits $n=8$, reduced Planck constant $\hbar=1$, Larmor frequency $\omega_{0}=1$, transverse field strength $\Omega=1$, and Heisenberg coupling constant $J=1$. For the Ising model, the anisotropy parameters are $\delta=1,\Delta=0$, corresponding to the standard Ising model, and the DMI strength is varied as $D\in\{0,5,10\}$ to capture baseline, intermediate, and strong interaction regimes. For the XXZ model, we set $\delta=0,\Delta=2$, representing anisotropy along the $z$-axis, and consider $D\in\{0,1.7\}$, where $D=1.7$ is chosen as a representative moderate value to illustrate the impact of finite DMI without reaching the saturation regime. In the following subsections, we compare the results for each Hamiltonian model across the different spin-chain configurations.
IV.1 Ising model
IV.1.1 Without DM interaction
@@
We first analyze the Ising model without DMI, as shown in Figs. LABEL:fig:erg_ising and LABEL:fig:power_ising. In Fig. LABEL:fig:erg_ising, the ergotropy of all three spin-chain configurations reaches the maximum value $2n\hbar\omega_{0}=16$ at $t=\pi/2$ for $\lambda=0$. This corresponds to full charging under the action of the charging Hamiltonian $H_{C}$ alone. As $\lambda$ increases, the maximum extractable energy decreases (orange curve, $\lambda=0.5$; green curve, $\lambda=1$), showing that the inclusion of $H_{B}$ suppresses energy storage across all configurations.
@@
Figure LABEL:fig:power_ising compares the charging power. The supercube configuration shows the highest peak power, followed by the closed chain and then the open chain. The peaks occur at $t<\pi/4$, indicating fast charging in all cases. For $\lambda>0$, the charging power decreases, with the divergence from the $\lambda=0$ curve becoming more pronounced at later times. This confirms that $H_{B}$ is detrimental both to the maximum stored energy and to the charging rate. Notably, the supercube retains relatively higher power even when $\lambda>0$, likely due to its richer interaction geometry.
IV.1.2 With DMI
@@
We next examine the effect of including DMI in the Ising model, shown in Figs. LABEL:fig:diff_d_erg_ising and LABEL:fig:diff_d_power_ising. Introducing DMI leads to a significant reduction in peak ergotropy compared with the $D=0$ case. Although increasing $D$ from 0 to 10 slightly raises ergotropy, the system still does not reach the fully charged state. For any $D$, increasing $\lambda$ further reduces the peak ergotropy, again reflecting the suppressive role of $H_{B}$.
@@
The effect on charging power depends on interaction geometry. For open boundary and supercube chains, larger $D$ enhances peak power and broadens the power curves, indicating higher charging rates sustained over longer times. The enhancement is particularly strong in the supercube configuration. By contrast, in the closed chain, increasing $D$ reduces the maximum power. For all cases, including $H_{B}$ ($\lambda>0$) visibly suppresses the peak power. Moreover, when $D$ greatly exceeds the Ising coupling $J$, the maximum power saturates, showing diminishing returns for very strong DMI.
@@ IV.2 XXZ model
As seen in the previous section, increasing DMI strength enhances the performance of the QB in the Ising model for open boundary and supercube configurations, while exhibiting a negative effect for the closed boundary configuration. We now study the effect of DMI on the three spin configurations of the QB in the XXZ model. We analyse ergotropy and charging power for $D=0$ and $1.7$, and also examine the influence of the battery Hamiltonian $H_{B}$ by varying the parameter $\lambda$, as shown in Figs. 4 and LABEL:fig:power_XXZ. The plots are presented for $t\in[0,3\pi]$. Results are discussed separately for each spin-chain configuration.
IV.2.1 Open boundary spin chain configuration
From the ergotropy plot in Fig. 4(a), we observe that introducing DMI does not enhance ergotropy. For $\lambda>0$, ergotropy is further reduced and exhibits irregular peaks, indicating that $H_{B}$ hinders energy storage. Unlike other topologies, the curves lack oscillatory behaviour, except for the case $D=0,\lambda=1$ which shows weak, low-amplitude oscillations. This suggests that DMI does not effectively facilitate energy transfer between the charging field and the battery for this topology.
@@ IV.2.1 Open boundary spin chain configuration
Figure LABEL:fig:power_XXZ(a) shows the charging power. Consistent with the ergotropy results, DMI provides little advantage. For $t\gtrsim 2$ up to $t\approx 6$, there is a slight increase in power for $D=1.7,\lambda=0$, but no significant enhancement in stability. For $\lambda>0$, peak powers are strongly suppressed, although nearly equal for $\lambda=0.5$ and $\lambda=1$, with DMI giving only a minor late-time improvement. Overall, the open chain gains limited benefit from DMI under the XXZ model.
IV.2.2 Closed boundary spin chain configuration
@@
In the closed boundary case [Fig. 4(b)], introducing DMI ($D=1.7$) leads to a marked increase in peak ergotropy, accompanied by sharp oscillations. However, the system discharges asymmetrically, i.e., after each discharge it does not return fully to the passive state but remains partially charged, leaving residual energy that cannot be extracted. This reduces cycle efficiency. Notably, the minima of the oscillations show an upward drift, indicating energy accumulation over repeated cycles. As $\lambda$ increases, peak ergotropy decreases, although the oscillatory nature of the curves is preserved. This indicates that DMI supports the periodic charging cycle of the QB system and including $H_{B}$ in the charging process reduces the maximum amount of energy that can be stored in the system. For $D=1.7$, ergotropy remains consistently higher than the $D=0$ case.
@@
Figure LABEL:fig:power_XXZ(b) shows that DMI significantly enhances the peak power for all $\lambda$. Increasing $\lambda$ reduces the peak values, but the DMI advantage persists, though weaker at later times. Thus, the closed boundary configuration benefits from DMI in the XXZ model, with the drawback of accumulating residual energy across cycles.
IV.2.3 Supercube spin chain configuration
@@
Figures 4(c) and LABEL:fig:power_XXZ(c) show that the supercube configuration demonstrates the most desirable features for a QB under the XXZ model. The ergotropy curve with DMI exhibits a clean sinusoidal charging–discharging cycle, periodically reaching the maximum value and returning to the passive state with negligible residual energy. This reset capability makes it well suited for rechargeable QB models and clocked quantum protocols. The sinusoidal behaviour remains stable over extended times, reflecting robust energy exchange between the battery and charging field facilitated by DMI [13]. For $\lambda>0$ at $D=1.7$, the cycle persists, though peak ergotropy decreases with increasing $\lambda$. Importantly, ergotropy remains well above the $D=0$ case throughout, reaffirming the positive role of DMI even in the presence of $H_{B}$ in the charging process.
@@
Looking at the power curves shown in figure LABEL:fig:power_XXZ(c), we note that supercube exhibits the highest peak among all three configurations for $D=1.7,\lambda=0$, slightly surpassing that for closed boundary spin chain. Power eventually decays over time but DMI leads to broader power curves. Overall, among the three configurations, the supercube consistently outperforms both open and closed chains in the XXZ model, mirroring its advantage in the Ising case.
@@
Across both Hamiltonian models, the supercube configuration consistently delivers superior performance, combining high ergotropy, strong charging power, and stable charging cycles. In the Ising model, DMI enhances charging power for open and supercube chains but degrades performance in the closed chain. In contrast, under the XXZ model, DMI significantly improves the closed and supercube chains, with the latter displaying clean sinusoidal cycles and negligible residual energy, making it particularly attractive for rechargeable QB applications. Thus, while the influence of DMI depends on both model and topology, the supercube architecture remains the most robust and efficient design. Motivated by the superior performance of the supercube configuration across both Ising and XXZ models, we now examine how variations in interaction parameters and structural modifications affect its charging behaviour.
@@
We study the effect of varying the DMI strength $D$ and the Heisenberg coupling constant $J$ on the charging behaviour of the supercube spin chain in XXZ model illustrated in figure LABEL:fig:erg_p_cube. Figure LABEL:fig-a:erg_vary_DM shows ergotropy for different values of $D$ keeping $J=1$ fixed. When $D\neq 1.7$, the optimal sinusoidal charging-discharging cycle disappears and the maximum ergotropy decreases, indicating that there exists an optimal DMI strength for sustaining high-efficiency periodic energy transfer between the battery and the charging field. Figure LABEL:fig-b:erg_vary_J shows ergotropy for different values of $J$ keeping $D=1.7$ fixed. The periodic charging cycle disappears for $J\neq 1$, and ergotropy of the system significantly diminishes, becoming nearly zero for $J>2$. These observations suggest that the combination $D=1.7$ and $J=1$ are optimal points for the supercube QB model under XXZ Hamiltonian. This behaviour further indicates the presence of a resonance-like condition, where the balance between exchange interaction ($J$) and spin–orbit coupling effects induced by DMI enables efficient cyclic energy transfer.

### v2

@@ II.1 Spin chain configurations
II.1.1 Open boundary spin chain configuration
In the open boundary configuration, the spins are arranged linearly under open boundary conditions, such that each spin interacts only with its nearest neighbor [Zheng2025]. The interaction Hamiltonian is given by
$H_{int}$ $=J\sum_{i=1}^{n-1}\Big[(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big]$ (11)
@@ II.1.1 Open boundary spin chain configuration
II.1.2 Closed boundary spin chain configuration
In the closed boundary configuration, the last spin interacts with the first spin, forming a closed loop. This implements periodic boundary conditions, ensuring that each spin has two neighbors, even at the boundaries [Zheng2025]. The interaction Hamiltonian is therefore similar to the open boundary case but extended to include the periodic connection between the first and last spins such that
$H_{int}$ $=J\sum_{i=1}^{n}\Big[(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+1}^{z}\Big]$ (12)
@@ II.1.2 Closed boundary spin chain configuration
II.1.3 Supercube spin chain configuration
The supercube configuration is inspired by the connectivity structure of quantum neural networks [QNN]. In this model, the system qubits are arranged in a one-dimensional array, but their interactions are mapped onto the structure of a cube, where the couplings between qubits are represented as edges. To further enrich the connectivity, additional couplings are introduced along the face diagonals and body diagonals. This extended interaction network allows us to study the effect of higher connectivity on the charging performance of the quantum battery. Consequently, the interaction Hamiltonian becomes more complex, incorporating nearest-neighbor, next-nearest-neighbor, and longer-range couplings, each accompanied by corresponding Dzyaloshinskii–Moriya (DM) interaction terms. The corresponding interaction Hamiltonian is
$H_{int}$ $=\Big\{J\sum_{i=1,i\in Z_{odd}}^{8}[(1+\delta)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+1}^{y}$ (13)
@@ II.1.3 Supercube spin chain configuration
$+\Big\{J\sum_{i=1,2,5,6}[(1+\delta)\sigma_{i}^{x}\sigma_{i+2}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+2}^{y}$
@@
$+\Big\{J\sum_{i=1}^{4}[(1+\delta)\sigma_{i}^{x}\sigma_{i+4}^{x}+(1-\delta)\sigma_{i}^{y}\sigma_{i+4}^{y}+\Delta\sigma_{i}^{z}\sigma_{i+4}^{z}]$
@@
(b) Ergotropy and Power for supercube spin chain configuration in XXZ model with varying interaction strengths
II.2 QB Performance Metrics
@@
II.2.1 Ergotropy
Ergotropy is defined as the maximum amount of work that can be extracted from a quantum battery [Allahverdyan_2004, Sone2021, Tirone2021, Touil_2022]. The concept relies on the notion of a passive state, from which no work can be extracted under unitary operations [Pusz1978, Lenard1978, Salvia2021distributionofmean, Mazzoncini2023].
If the battery Hamiltonian can be written as $H_{B}=\sum_{i}{\epsilon_{i}|\epsilon_{i}\rangle\langle\epsilon_{i}|}$ where $\epsilon_{i}$ represents the eigenenergies of $H_{B}$ and $|\epsilon_{i}\rangle$ are the corresponding eigenstates ordered by $\epsilon_{i}<\epsilon_{i+1}$, and the density matrix of the system be $\rho(t)=\sum_{i}{r_{i}|r_{i}\rangle\langle r_{i}|}$ such that $r_{i}\geq r_{i+1}$ where $r_{i}$ represent the eigenvalues of $\rho(t)$ and $|r_{i}\rangle$ are the corresponding eigenvectors, then passive state of the system can be defined as $\rho_{p}=\sum_{i}{r_{i}}|\epsilon_{i}\rangle\langle\epsilon_{i}|$ [Francica2020]. Thus ergotropy will be
$\zeta=\text{Tr}[H_{B}(\rho(t)-\rho_{p})].$ (14)
For pure states, the ground state of the system is passive ($\rho_{p}={r_{0}}|\epsilon_{0}\rangle\langle\epsilon_{0}|$, $r_{0}=1$) [Passive_state]. Accordingly, the ergotropy is given by the difference in energy between the evolved state $|\psi(t)\rangle=U|\psi(0)\rangle$ and the ground state $|\psi(0)\rangle$, i.e.
$\zeta=\langle\psi(t)|H_{B}|\psi(t)\rangle-\epsilon_{0},$ (15)
where, $|\psi(0)\rangle=(|\downarrow\rangle)^{\otimes n}$ is the initial state of the system, which is also the ground state of $H_{B}$, and $\epsilon_{0}$ denotes passive state energy, i.e., the lowest eigenvalue of $H_{B}$ given by $\langle\psi(0)|H_{B}|\psi(0)\rangle$. $H_{B}$ denotes the internal Hamiltonian of the battery as defined in Eq. (5).
(c) Adding connections corresponding to body diagonal and face diagonal
II.2.2 Charging Power
@@
III Results
In this section we present the results for the eight-qubit quantum battery under different spin-chain configurations and Hamiltonian models. Our analysis focuses on the Ising and XXZ models, and compares their performance in terms of ergotropy and charging power. We also examine the influence of the battery Hamiltonian $H_{B}$ on the charging process by varying the parameter $\lambda\in[0,1]$. In addition, we extend our analysis of QB architecture in XXZ model to varying system sizes, to explore how number of qubits influence charging performance. The population dynamics plots for all three geometries in Ising model and XXZ model for $D=0,1.7$ when $\lambda=0$ are included in the Appendix A. While the main text discusses the qualitative trends and their implications for energy transfer and ergotropy, the detailed population evolution offers an additional layer of transparency regarding how individual spin levels redistribute during the charging process. These plots help verify that the observed charging behavior is not an artefact of specific initial conditions or restricted parameter choices, and they serve as a diagnostic tool to distinguish whether features such as enhanced charging arise from interaction-driven dynamics rather than entanglement effects. Including these results in the Appendix allows one to examine the full dynamical behavior without interrupting the flow of the main discussion, while still ensuring that all supporting analyzes remain accessible.
@@ III Results
Unless stated otherwise, the values of the parameters used in this study are as follows: number of qubits $n=8$, reduced Planck constant $\hbar=1$, Larmor frequency $\omega_{0}=1$, transverse field strength $\Omega=1$, and Heisenberg coupling constant $J=1$. For the Ising model, the anisotropy parameters are $\delta=1,\Delta=0$, corresponding to the standard Ising model, and the DMI strength is varied as $D\in\{0,1,2\}$ to examine the effect of increasing interaction strength. For the XXZ model, we set $\delta=0,\Delta=2$, representing anisotropy along the $z$-axis, and consider $D\in\{0,1.7\}$, where $D=1.7$ is chosen as a representative moderate value to illustrate the impact of finite DMI without reaching the saturation regime. In the following subsections, we compare the results for each Hamiltonian model across the different spin-chain configurations.
III.1 Ising model
III.1.1 Without DM interaction
@@
We first analyze the Ising model without DMI, as shown in Figs. 4(a) and 4(b). We study the effect of the battery Hamiltonian $H_{B}$ on the ergotropy of the system and compare the charging power across the three configurations, evolving each system until it reaches its maximum ergotropy value =16=16 (i.e., at $t=\pi/2$). In Fig. 4(a), the ergotropy of all three spin-chain configurations reaches the maximum value $2n\hbar\omega_{0}=16$ at $t=\pi/2$ for $\lambda=0$. This corresponds to full charging under the action of the charging Hamiltonian $H_{C}$ alone. As $\lambda$ increases, the maximum extractable energy decreases (orange curve, $\lambda=0.5$; green curve, $\lambda=1$), showing that the inclusion of $H_{B}$ suppresses energy storage across all configurations.
@@
Figure 4(b) compares the charging power. The supercube configuration shows the highest peak power, followed by the closed chain and then the open chain. The peaks occur at $t<\pi/4$, indicating fast charging in all cases. For $\lambda>0$, the charging power decreases, with the divergence from the $\lambda=0$ curve becoming more pronounced at later times. This confirms that $H_{B}$ is detrimental both to the maximum stored energy and to the charging rate. Notably, the supercube retains relatively higher power even when $\lambda>0$, likely due to its richer interaction geometry.
III.1.2 With DMI
@@
We next examine the effect of including DMI in the Ising model and analyze the changes in peak power and ergotropy curves for $D=0,1,2$ when the system is evolved until ergotropy attains its maximum value of $16$ for $D=0$, shown in Figs. 4 and 7(a). Introducing DMI leads to a significant reduction in peak ergotropy compared with the $D=0$ case. For any $D$, increasing $\lambda$ further reduces the peak ergotropy, again reflecting the suppressive role of $H_{B}$.
@@
The effect on charging power depends on interaction geometry. For supercube chain, larger $D$ enhances peak power, indicating higher charging rates. While in open boundary spin chain, increasing $D$ from 0 to 2 does not show any significant improvement in maximum charging power. The three curves ($D=0,1,2$) overlap until $t\approx\pi/4$ and diverge slightly at later times. Remarkably, when $D$ is increased beyond 2, peak power for the system also increases. By contrast, in the closed chain, increasing $D$ reduces the maximum power. For all cases, including $H_{B}$ ($\lambda>0$) visibly suppresses the peak power. Moreover, when $D$ greatly exceeds the Ising coupling $J$, the maximum power saturates, showing diminishing returns for very strong DMI [zhang2024quantum].
@@ III.2 XXZ model
As seen in the previous section, increasing DMI strength enhances the performance of the QB in the Ising model for supercube configuration, while exhibiting a negative effect for the closed boundary configuration. We now study the effect of DMI on the three spin configurations of the QB in the XXZ model. We analyze ergotropy and charging power for $D=0$ and $1.7$, and also examine the influence of the battery Hamiltonian $H_{B}$ by varying the parameter $\lambda$, as shown in Figs. 7 and 10(a). The plots are presented for $t\in[0,3\pi]$ as it captures at least one entire oscillation cycle. For reference, the maximum ergotropy for our 8 qubit system is 16. And our focus is on the periodicity of ergotropy curve as it indicates the ability of the system to fully charge and then reset to 0. Results are discussed separately for each spin-chain configuration.
III.2.1 Open boundary spin chain configuration
From the ergotropy plot in Fig. 7(a), we observe that the curves lack regular oscillatory behavior. The system never reaches its fully charged state, hence ergotropy is always less than the maximum value of $16$. Even upon introducing DMI and the battery Hamiltonian into the system, the value of ergotropy fluctuates within a certain range without reaching higher amplitudes.
@@ III.2.1 Open boundary spin chain configuration
Figure 10(a)(a) shows the corresponding charging power. Consistent with the ergotropy results, DMI provides little advantage. For intermediate times $t\gtrsim 2$ up to $t\approx 6$, there is a slight increase in power for $D=1.7,\lambda=0$, but no significant enhancement in stability. For $\lambda>0$, we observe an increase in peak power which increases as $\lambda$ increases, with DMI suppressing the peak value initially and later converging to the $D=0$ case. Overall, the open chain gains limited benefit from DMI under the XXZ model.
III.2.2 Closed boundary spin chain configuration
@@
In the closed boundary case [Fig. 7(b)], introducing DMI ($D=1.7$) leads to a marked increase in peak ergotropy, accompanied by sharp oscillations. However, the system discharges asymmetrically, i.e., after each discharge it does not return fully to the passive state but remains partially charged, leaving residual energy that cannot be extracted. This reduces cycle efficiency. Notably, the minima of the oscillations show an upward drift, indicating energy accumulation over repeated cycles. When $\lambda=0.5$, peak ergotropy increases but when $\lambda=1$, there is a drop in peak value. For $D=1.7$, ergotropy remains consistently higher than the $D=0$ case for $\lambda=0,1$, but for $\lambda=0.5$ both the curves initially follow the same trend, but diverge significantly afterwards.
@@
Figure 10(a)(b) shows that DMI significantly enhances the peak power for $\lambda=0$ and similarly for $\lambda=0.5$ DMI results in a slight increase in peak value. However, for $\lambda=1$, DMI suppresses maximum power. Increasing $\lambda$ increases the increases the maximum power for $D=0$ case, but a slight decrease is observed for $D=1.7$ when $\lambda=1$. Thus, the closed boundary configuration benefits from DMI in the XXZ model, with the drawback of accumulating residual energy across cycles.
(d) Structures (ac)(a-c) and Ergotropy time evolution (df)(d-f) for symmetric Platonic geometries: (a) tetrahedron $(n=4)$, (b) octahedron $(n=6)$, and (c) icosahedron $(n=12)$. All cases exhibit perfect sinusoidal charging–discharging cycles at their respective optimal DMI strengths, demonstrating that highly symmetric connectivities preserve ideal QB operation across different system sizes.
III.2.3 Supercube spin chain configuration
@@
Figures 7(c) and 10(a)(c) show that the supercube configuration demonstrates the most desirable features for a QB under the XXZ model. The ergotropy curve with DMI exhibits a clean sinusoidal charging–discharging cycle, periodically reaching close to the maximum value and returning to the passive state with negligible residual energy. This reset capability makes it well suited for rechargeable QB models and clocked quantum protocols. The sinusoidal behavior remains stable over extended times, reflecting robust energy exchange between the battery and charging field facilitated by DMI [energy_transfer]. For $\lambda>0$ at $D=1.7$, the cycle persists, though peak ergotropy decreases with increasing $\lambda$. Importantly, ergotropy remains well above the $D=0$ case throughout except when $\lambda=1$, showing that inclusion of the battery Hamiltonian plays a detrimental role in the charging dynamics.
@@
Looking at the power curves shown in figure 10(a)(c), we note that supercube exhibits the highest peak among all three configurations for $D=1.7$ when $\lambda=0$, slightly surpassing that for closed boundary spin chain. Power eventually decays over time but DMI leads to broader power curves except when the battery Hamiltonian is included in the charging dynamics. Overall, among the three configurations, the supercube consistently outperforms both open and closed chains in the XXZ model for $\lambda=0$, mirroring its advantage in the Ising case.
@@
Across all the Hamiltonian models, the supercube configuration consistently delivers superior performance, combining high ergotropy, strong charging power, and stable charging cycles. In the Ising model, DMI enhances charging power for open and supercube chains but degrades performance in the closed chain. In contrast, under the XXZ model, DMI significantly improves the closed and supercube chains, with the latter displaying clean sinusoidal cycles and negligible residual energy, making it particularly attractive for rechargeable QB applications. Thus, while the influence of DMI depends on both model and topology, the supercube architecture remains the most robust and efficient design. Motivated by the superior performance of the supercube configuration across both Ising and XXZ models, we now examine how variations in interaction parameters and structural modifications affect its charging behavior.
@@
We study the effect of varying the DMI strength $D$ and the Heisenberg coupling constant $J$ on the charging behavior of the supercube spin chain in XXZ model illustrated in figure 10(b). Figure LABEL:fig-a:erg_vary_DM shows ergotropy for different values of $D$ keeping $J=1$ fixed. When $D\neq 1.7$, the optimal sinusoidal charging-discharging cycle disappears and the maximum ergotropy decreases, indicating that there exists an optimal DMI strength for sustaining high-efficiency periodic energy transfer between the battery and the charging field. Figure LABEL:fig-b:erg_vary_J shows ergotropy for different values of $J$ keeping $D=1.7$ fixed. The periodic charging cycle disappears for $J\neq 1$, and ergotropy of the system significantly diminishes, becoming nearly zero for $J>2$. These observations suggest that the combination $D=1.7$ and $J=1$ are optimal points for the supercube QB model under XXZ Hamiltonian. This behavior further indicates the presence of a resonance-like condition, where the balance between exchange interaction ($J$) and spin–orbit coupling effects induced by DMI enables efficient cyclic energy transfer.